OpenStudy (anonymous):

Determine if 3 * 4^1/6 is the simplest form of 2916^1/6. If not, tell why not and write the expression in simplest form.

OpenStudy (anonymous):

4^(1/6) = 2^(2/6)

OpenStudy (anonymous):

= 2^(1/3)